OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2 + 10x + 22 =

OpenStudy (anonymous):

\[x^2+10x=-22\] \[(x+5)^2=-22+5^2=-22+25=3\] \[x+5=\pm\sqrt{3}\] \[x=-5\pm\sqrt{3}\]